Two circular coils are made from a uniform wire the ratio of radii of circular coils are 2:3 and no. of turns is 3:4. If they are connected in parallel across a battery. Find ratio of magnetic inductions at their centres.

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

The magnetic field at the center of each coill is given by-

B=\dfrac{\mu_0Ni}{2R}
\implies \dfrac{B_2}{B_1}=\dfrac{N_2}{N_1}\times \dfrac{I_2}{I_1}\times\dfrac{R_1}{R_2}
Also the ratio of resistances will be\dfrac{r_2}{r_1}=\dfrac{l_2}{l_1}=\dfrac{N_2(2\pi R_2)}{N_1(2\pi R_1)},
\implies \dfrac{I_2}{I_1}=\dfrac{r_2}{r_1}=\dfrac{N_1R_1}{N_2R_2}
Use these ratios to find the answer.